# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ Appendix A Proof of Proposition 1
Appendix B Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
@@
We start by recalling the definition of subdifferentiability in Wasserstein spaces, see [4, Definition 10.1.1]. We denote by $\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ the space of probability distributions in $\mathcal{P}_{2}(\mathbb{R}^{d})$ which are absolutely continuous w.r.t the Lebesgue measure and equip the space $\mathrm{L}^{2}(\mathbb{R}^{d},\pi):=\{f:\mathbb{R}^{d}\to\mathbb{R}^{d};\pi(\|f\|^{2})<\infty\}$ with the norm $\left\|f\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}:=\pi(\|f\|^{2})$.
@@
Under Assumption 1, the strong subdifferential of $\mathcal{G}^{\eta}$ is given by
@@
Let $\mathcal{L}:\ \mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{p}\to\left[0,+\infty\right)$ be given for any $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{d}$ by
$\mathcal{L}(\pi,x)=\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\varphi(x)$
and denote by $\xi_{1}(x,z):=\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)$ so that $\nabla\mathcal{L}(\pi,x)=\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z)$ for all fixed $\pi\in\mathcal{P}(\mathbb{R}^{d})$, where Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]) guarantees that we can swap the derivative w.r.t. $x$ with the integral w.r.t. $z$ since $\left\|\nabla_{1}\mathrm{k}\right\|\leq\mathtt{M}$ and $\mathrm{k}$ is bounded and thus integrable.
Further, define $\xi_{2}(x,z):=\lambda\nabla_{2}\mathrm{k}(x,z)$, let $g:\ \left[0,+\infty\right)\to\mathbb{R}$ be given for any $t\geq 0$ by $g(t)=-\log(t+\eta)$ and consider
@@
For the second group of terms in (21), we have, using Assumption 1 and the fact that $\mathcal{L}(\pi,x)$ is assumed to be positive for all $\pi\in\mathcal{P}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}$, we have, using the spectral norm induced by the Euclidean norm of vectors,
@@
$\leq$ $\frac{((\lambda+1)\mathtt{M}+\eta)(\lambda+1)\mathtt{M}+(\lambda+1)^{2}\mathtt{M}^{2}}{\eta^{2}}:=\mathtt{C}_{0},$
where $\nabla\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)=\int\limits_{\mathbb{R}^{d}}\nabla^{2}_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)$ using Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]) since $\left\|\nabla_{1}^{2}\mathrm{k}\right\|\leq\mathtt{M}$ and $\xi_{1}$ is bounded and thus integrable. Using that $\left\|\nabla^{2}g(\mathcal{L}(\pi,x))\right\|\leq\mathtt{C}_{0}$ we get that for any $x_{1},x_{2}\in\mathbb{R}^{d}$,
@@
and [4, Proposition 10.4.2] gives that the limit as $n\to\infty$ of
$\frac{\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}(t_{n\#}\pi-\pi)(x)-\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,x))\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$ (24)
is non-negative. Combining (21)–(24) we obtain
@@
$\;\phantom{\geq}-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x).$

### v2

@@ A.2 Proof of Proposition 1
Appendix B Convergence of Minimizers
B.1 Basics on $\Gamma$-convergence
We start this section by recalling some basic facts on $\Gamma$-convergence. We refer to [19] for a more complete introduction to $\Gamma$-convergence.
First, we recall that a function $f:\ \mathsf{X}\to\mathbb{R}$ (where $\mathsf{X}$ is a metric space) is coercive if for any $t\in\mathbb{R}$, $f^{-1}(\left(-\infty,t\right])$ is relatively compact. This definition can be extended to the case where $\mathsf{X}$ is only a topological space, see [19, Definition 1.12].
Definition 1.
Let $F=\{f_{\alpha}:\,\alpha\in\mathsf{A}\}$ where $\mathsf{A}$ is a topological space and for any $\alpha\in\mathsf{A}$, $f_{\alpha}:\mathsf{X}\to\mathbb{R}\cup\{+\infty\}$ with $\mathsf{X}$ a metric space. We say that $F$ is a $\Gamma$-system if the following hold:
  1. (a)
    For any $x\in\mathsf{X}$, $\alpha^{\star}\in\mathsf{A}$, $(\alpha_{n})_{n\in\mathbb{N}}\in\mathsf{A}^{\mathbb{N}}$ with $\lim_{n\to+\infty}\alpha_{n}=\alpha^{\star}$ and $(x_{n})_{n\in\mathbb{N}}\in\mathsf{X}^{\mathbb{N}}$ with $\lim_{n\to+\infty}x_{n}=x$ we have $\liminf_{n\to+\infty}f_{\alpha_{n}}(x_{n})\geq f_{\alpha^{\star}}(x)$.
  2. (b)
    For any $x\in\mathsf{X}$, $\alpha^{\star}\in\mathsf{A}$, $(\alpha_{n})_{n\in\mathbb{N}}\in\mathsf{A}^{\mathbb{N}}$ with $\lim_{n\to+\infty}\alpha_{n}=\alpha^{\star}$ there exists $(x_{n})_{n\in\mathbb{N}}\in\mathsf{X}^{\mathbb{N}}$ with $\lim_{n\to+\infty}x_{n}=x$ such that $\lim_{n\to+\infty}f_{\alpha_{n}}(x_{n})=f_{\alpha^{\star}}(x)$.
Proposition 8 (Corollary 7.20, [19].).
Let $F=\{f_{\alpha}:\,\alpha\in\mathsf{A}\}$ where $\mathsf{A}$ is a topological space and for any $\alpha\in\mathsf{A}$, $f_{\alpha}:\mathsf{X}\to\mathbb{R}\cup\{+\infty\}$ with $\mathsf{X}$ a metric space. Assume that $F$ is a $\Gamma$-system and is such that for any $\alpha\in\mathsf{A}$, $f_{\alpha}$ admits a unique minimizer. Then for any $(\alpha_{n})_{n\in\mathbb{N}}\in\mathsf{A}^{\mathbb{N}}$ and $(x_{n})_{n\in\mathbb{N}}\in\mathsf{X}^{\mathbb{N}}$ such that for any $n\in\mathbb{N}$, $x_{n}$ is the minimizer of $f_{\alpha_{n}}$ and $\lim_{n\to+\infty}\alpha_{n}=\alpha^{\star}$, if there exists $x^{\star}\in\mathsf{X}$ such that $\lim_{n\to+\infty}x_{n}=x^{\star}$ then $x^{\star}$ is a minimizer of $f_{\alpha^{\star}}$.
B.2 Proof of Proposition 2
We start by showing an auxiliary result.
Lemma 2.
Under Assumptions 1 and 2, for any $\alpha>0$, $\eta\geq 0$, $\pi_{\alpha,\eta}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$.
Proof.
Take $\alpha>0$ and $\eta\geq 0$. Since by Proposition 1, $\mathcal{F}_{\alpha}^{\eta}$ is proper, $\mathcal{F}_{\alpha}^{\eta}(\pi_{\alpha,\eta}^{\star})<+\infty$ and therefore, $\textup{KL}\left(\pi_{\alpha,\eta}^{\star}\middle|\pi_{0}\right)<+\infty$. Hence, $\pi_{\alpha,\eta}^{\star}$ admits a density w.r.t to $\pi_{0}$. Since $\pi_{0}$ is dominated by the Lebesgue measure, we get that $\pi_{\alpha,\eta}^{\star}$ admits a density with respect to the Lebesgue measure.
In addition, let $\pi_{1}\in\mathcal{P}(\mathbb{R}^{d})$ be such that for any $x\in\mathbb{R}^{d}$ we have for some $\tau>0$
$(\mathrm{d}\pi_{1}/\mathrm{d}\pi_{0})(x)=f(x)/\pi_{0}[f],\qquad f(x)=\exp[(\tau/2)\left\|x\right\|^{2}].$
Since $\textup{KL}\left(\pi_{\alpha,\eta}^{\star}\middle|\pi_{0}\right)<+\infty$ we have $\int_{\mathbb{R}^{d}}|\log((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{0})(x))|\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)<+\infty$ and
$\int_{\mathbb{R}^{d}}\log((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathbbm{1}_{[1,+\infty)}((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)$ (22)
$\qquad\quad=\int_{\mathbb{R}^{d}}\left[\log((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{0})(x))-(\tau/2)\left\|x\right\|^{2}+\log(\pi_{0}[f])\right]$
$\qquad\qquad\times\mathbbm{1}_{[1,+\infty)}((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)$
$\qquad\quad\leq\int_{\mathbb{R}^{d}}\left[|\log((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{0})(x))|+\log(\pi_{0}[f])\right]$
$\qquad\qquad\times\mathbbm{1}_{[1,+\infty)}((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)<+\infty.$
Since $t\mapsto t\log(t)$ is bounded on $[0,1]$ we have that
$\int_{\mathbb{R}^{d}}\log((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathbbm{1}_{[0,1]}((\mathrm{d}\pi_{\alpha,\eta}^{\star}/\mathrm{d}\pi_{1})(x))\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)<+\infty.$ (23)
Combining (23) and (22) we get that $\textup{KL}\left(\pi_{\alpha,\eta}^{\star}\middle|\pi_{1}\right)<+\infty$. Therefore, using [17, Equation 2.6], we get that
$(\tau/2)\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)-\log(\pi_{0}[f])$
$\qquad=\int_{\mathbb{R}^{d}}\log((\mathrm{d}\pi_{1}/\mathrm{d}\pi_{0})(x))\mathrm{d}\pi_{\alpha,\eta}^{\star}(x)=\textup{KL}\left(\pi_{\alpha,\eta}^{\star}\middle|\pi_{0}\right)-\textup{KL}\left(\pi_{\alpha,\eta}^{\star}\middle|\pi_{1}\right)<+\infty.$
Therefore, $\pi_{\alpha,\eta}^{\star}\in\mathcal{P}_{2}(\mathbb{R}^{d})$.
We can now show that the results in Proposition 2 hold.
  1. 1.
    By Lemma 2, $\pi_{\alpha,\eta_{n}}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. Consider for fixed $\alpha$ the family $\{\mathcal{F}_{\alpha}^{\eta}:\eta\geq 0\}$. Let $(\eta_{n})_{n\in\mathbb{N}^{\star}}\in[0,+\infty)^{\mathbb{N}^{\star}}$ be such that $\underset{n\to+\infty}{\lim}\eta_{n}=0$ and $\underset{n\to+\infty}{\lim}\pi_{\alpha,\eta_{n}}^{\star}=\pi_{\alpha}^{\star}$ with $\pi^{\star}_{\alpha}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. For any $(\pi_{n})_{n\in\mathbb{N}^{\star}}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})^{\mathbb{N}^{\star}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$ we have using Fatou’s lemma
    $\liminf_{n\to+\infty}\mathcal{F}_{\alpha}^{\eta_{n}}(\pi_{n})\geq\mathcal{F}_{\alpha}(\pi).$
    This guarantees condition (a) in Definition 1 holds.
    To check condition (b) in Definition 1 consider w.l.o.g. $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ which assigns positive mass to $1,1]^{d}$. For any $n\in\mathbb{N}^{\star}$, we denote $\pi_{n}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ with density w.r.t. the Lebesgue measure given for any $x\in\mathbb{R}^{d}$ by
    $\pi_{n}(x)=\frac{\pi(x)\mathds{1}_{[-n,n]^{d}}(x)}{Z_{n}},\qquad\textrm{where }Z_{n}:=\int_{[-n,n]^{d}}\pi(x)\mathrm{d}x.$ (24)
    Therefore, we have $\lim_{n\to+\infty}\mathbf{W}_{2}(\pi_{n},\pi)^{2}=0$ as $\pi_{n}\to\pi$ weakly and $\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi_{n}(x)\mathrm{d}x\to\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\pi(x)\mathrm{d}x$. We also observe that by construction $0<Z_{1}\leq Z_{2}\leq\dots\leq 1$ and $|\pi_{n}(x)|\leq\pi(x)/Z_{1}$.
    We now show that $\lim_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})=\mathcal{F}(\pi)$. Consider the function $x\mapsto\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})$. For any $n\in\mathbb{N}^{\star}$ and $x\in\mathbb{R}^{d}$, $\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log(2\lambda\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})$ and thus
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\leq\log(2\lambda\mathtt{M}+\sup_{n\in\mathbb{N}}\eta_{n})\frac{\pi(x)}{Z_{1}}.$ (25)
    Define $\Phi:\ \mathbb{R}^{d}\to\mathbb{R}$, $\Phi:x\mapsto\left\|x\right\|^{2}$. Using Jensen’s inequality and the fact that under Assumption 2 there exists $C_{2}\geq 0$ such that for any $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $\exp[-\Phi(x)-\Phi(y)]/\mathrm{k}(x,y)\leq C_{2}\exp[C_{2}]$, we have for any $y\in\mathbb{R}^{d}$ and $n\in\mathbb{N}^{\star}$ that
    $\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})$ $\geq\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)])$
    $\geq\lambda\left(-\Phi(x)-\sup_{n\in\mathbb{N}^{\star}}\pi_{n}[\Phi]-C_{2}-\log C_{2}\right),$
    where we used the fact that $\varphi(x)\geq 0$ under Assumption 1. Since, $(\pi_{n})_{n\in\mathbb{N}^{\star}}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$ [4, 7.1.5], there exists $C\geq 0$ such that for any $n\in\mathbb{N}^{\star}$, $\pi_{n}[\Phi]\leq C$, see [56, Definition 6.8]. Hence,
    $\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})\geq\lambda\left(-\Phi(x)-C-C_{2}-\log C_{2}\right)\frac{\pi(x)}{Z_{1}}.$ (26)
    Combining (25)–(26) and the fact that $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, the function $x\mapsto\pi_{n}(x)\log(\varphi(x)+\lambda\pi_{n}[\mathrm{k}(x,\cdot)]+\eta_{n})$ is bounded by an integrable function.
    Consider now $x\mapsto\pi_{n}(x)\log\pi_{n}(x)$ and observe that
    $\pi_{n}(x)\log\pi_{n}(x)=\pi_{n}(x)\log\pi(x)+\log(1/Z_{n})\pi_{n}(x)$
    As $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$, its entropy $\mathrm{H}(\pi)$ is finite and from the definition of the Lebesgue integral we have that $\int_{\mathbb{R}^{d}}\pi(x)|\log\pi(x)|\mathrm{d}x<\infty$.
    We can further bound
    $|\pi_{n}(x)\log\pi_{n}(x)|\leq$ $\pi_{n}(x)|\log\pi(x)|+\log(1/Z_{n})\pi_{n}(x)$
    $\leq$ $\pi_{n}(x)(|\log\pi(x)|+\log(1/Z_{1}))$
    $\leq$ $\pi(x)(|\log\pi(x)|+\log(1/Z_{1})|)/Z_{1}$ (27)
    which is an integrable function whose integral is bounded.
    Hence, noting that $\pi_{n}(x)\log\pi_{n}(x)$ converges pointwise to $\pi(x)\log\pi(x)$, we have that $\int_{\mathbb{R}^{d}}\pi_{n}(x)\log\pi_{n}(x)\mathrm{d}x\to\int_{\mathbb{R}^{d}}\pi(x)\log\pi(x)\mathrm{d}x$ by the Dominated Convergence Theorem, and that $\lim_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})=\mathcal{F}(\pi)$.
    In what follows, we show that $\lim_{n\to+\infty}\textup{KL}\left(\pi_{n}\middle|\pi_{0}\right)=\textup{KL}\left(\pi\middle|\pi_{0}\right)$. Using Assumption 2 there exists $C_{1}\geq 0$ and $\tau>0$ such that for any $x\in\mathbb{R}^{d}$, $|U(x)|\leq C_{1}+\tau\left\|x\right\|^{2}$. Since $\pi_{n}$ is upper bounded by an integrable function, so is $x\mapsto U(x)\pi_{n}(x)$ and the dominated convergence theorem guarantees $\int_{\mathbb{R}^{d}}U(x)\pi_{n}(x)\mathrm{d}x\to\int_{\mathbb{R}^{d}}U(x)\pi(x)\mathrm{d}x$ as $n\to\infty$. Then using the boundedness of $x\mapsto\pi_{n}(x)\log\pi_{n}(x)$ established above and the dominated convergence theorem we immediately have $\lim_{n\to+\infty}\textup{KL}\left(\pi_{n}\middle|\pi_{0}\right)=\textup{KL}\left(\pi\middle|\pi_{0}\right)$.
    Therefore, we have that $\mathcal{F}_{\alpha}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1. Using Proposition 8, we get that $\pi_{\alpha}^{\star}$ is a minimizer of $\mathcal{F}_{\alpha}$.
  2. 2.
    By Lemma 2, $\pi_{\alpha_{n},\eta_{n}}^{\star}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. Consider $\{\mathcal{F}_{\alpha}^{\eta}:\alpha,\eta\geq 0\}$. Let $(\alpha_{n})_{n\in\mathbb{N}}\in(0,+\infty)^{\mathbb{N}},(\eta_{n})_{n\in\mathbb{N}}\in[0,+\infty)^{\mathbb{N}}$ be such that $\underset{n\to+\infty}{\lim}\alpha_{n}=0$, $\underset{n\to+\infty}{\lim}\eta_{n}=0$. For any $(\pi_{n})_{n\in\mathbb{N}}\in\mathcal{P}^{ac}(\mathbb{R}^{d})^{\mathbb{N}}$ such that $\lim_{n\to+\infty}\pi_{n}=\pi$ we have using the monotonicity in $\alpha$ of $\mathcal{F}_{\alpha}^{\eta}$ and Fatou’s lemma
    $\liminf_{n\to+\infty}\mathcal{F}_{\alpha_{n}}^{\eta_{n}}(\pi_{n})\geq\liminf_{n\to+\infty}\mathcal{F}^{\eta_{n}}(\pi_{n})\geq\mathcal{F}(\pi).$
    This guarantees condition (a) in Definition 1 holds.
    To check condition (b) in Definition 1 consider for $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ and for any $n\in\mathbb{N}$, $\pi_{n}\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ defined as in (24). The proof of point 1 shows that $\mathcal{F}^{\eta_{n}}(\pi_{n})\to\mathcal{F}(\pi)$ as $n\to\infty$.
    We now show that $\lim_{n\to+\infty}\alpha_{n}\textup{KL}\left(\pi_{n}\middle|\pi_{0}\right)=0$. Using Assumption 2 there exists $C_{1}\geq 0$ and $\tau>0$ such that for any $x\in\mathbb{R}^{d}$, $|U(x)|\leq C_{1}+\tau\left\|x\right\|^{2}$. Since $|\pi_{n}(x)|\leq\pi(x)/Z_{1}$ we have
    $\int_{\mathbb{R}^{d}}|U(x)|\pi_{n}(x)\mathrm{d}x$ $\leq\frac{C_{1}}{Z_{1}}+\tau\int_{\mathbb{R}^{d}}\left\|x\right\|^{2}\frac{\pi(x)}{Z_{1}}\mathrm{d}x<+\infty$ (28)
    since $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$. Using (27) we also have that
    $|\pi_{n}(x)\log\pi_{n}(x)|$ $\leq\pi(x)\frac{|\log\pi(x)|+\log 1/Z_{1}}{Z_{1}}:=g(x)$ (29)
    with $g(x)$ integrable since $\pi\in\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ implies $\mathrm{H}(\pi)<\infty$.
    Combining (28) and (29) we have $\lim_{n\to+\infty}\alpha_{n}\textup{KL}\left(\pi_{n}\middle|\pi_{0}\right)=0$.
    Therefore, we have that $\mathcal{F}(\pi)=\lim_{n\to+\infty}\mathcal{F}_{\alpha_{n}}^{\eta_{n}}(\pi_{n})$, and $\{\mathcal{F}_{\alpha}^{\eta}:\,\alpha,\eta\geq 0\}$ is a $\Gamma$-system. In addition, for any $\alpha,\eta>0$, $\mathcal{F}_{\alpha}^{\eta}$ admits a unique minimizer $\pi_{\alpha,\eta}^{\star}$ by Proposition 1. Using Proposition 8, we get that $\pi^{\star}$ is a minimizer of $\mathcal{F}$.
Appendix C Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
@@
We start by recalling the definition of subdifferentiability in Wasserstein spaces, see [4, Definition 10.1.1]. We denote by $\mathcal{P}_{2}^{ac}(\mathbb{R}^{d})$ the space of probability measures in $\mathcal{P}_{2}(\mathbb{R}^{d})$ which are absolutely continuous w.r.t the Lebesgue measure and equip the space $\mathrm{L}^{2}(\mathbb{R}^{d},\pi):=\{f:\mathbb{R}^{d}\to\mathbb{R}^{d};\pi(\|f\|^{2})<\infty\}$ with the norm $\left\|f\right\|^{2}_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}:=\pi(\|f\|^{2})$.
@@
Under Assumption 1, $\mathcal{G}^{\eta}$ has strong subdifferential
@@
Let $\mathcal{L}:\ \mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\left[0,+\infty\right)$ be given for any $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{d}$ by $\mathcal{L}(\pi,x)=\lambda\int\limits_{\mathbb{R}^{d}}\mathrm{k}(x,z)\mathrm{d}\pi(z)+\varphi(x)$, and denote by $\xi_{1}(x,z):=\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)$ so that $\nabla\mathcal{L}(\pi,x)=\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z)$ for all fixed $\pi\in\mathcal{P}(\mathbb{R}^{d})$, where Leibniz integral rule (e.g. [8, Theorem 16.8]) guarantees that we can swap the derivative w.r.t. $x$ with the integral w.r.t. $z$ since $\left\|\nabla_{1}\mathrm{k}\right\|\leq(1+\lambda)\mathtt{M}$ and $\mathrm{k}$ is bounded and thus integrable.
[…]
@@
For the second group of terms in (30), using Assumption 1 and that $\mathcal{L}(\pi,x)$ is positive for all $\pi\in\mathcal{P}(\mathbb{R}^{d}),x\in\mathbb{R}^{d}$, we have, using the spectral norm induced by the Euclidean norm,
@@
$\leq$ $\frac{((\lambda+1)\mathtt{M}+\eta)(\lambda+1)\mathtt{M}+(\lambda+1)^{2}\mathtt{M}^{2}}{\eta^{2}}=:\mathtt{C}_{0},$
where $\nabla\int\limits_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)=\int\limits_{\mathbb{R}^{d}}\nabla^{2}_{1}\mathrm{k}(x,z)\mathrm{d}\pi(z)$ using Leibniz integral rule for differentiation under the integral sign (e.g. [8, Theorem 16.8]) since $\left\|\nabla_{1}^{2}\mathrm{k}\right\|\leq\mathtt{M}$ and $\xi_{1}$ is bounded and thus integrable. Using that $\left\|\nabla^{2}g(\mathcal{L}(\pi,x))\right\|\leq\mathtt{C}_{0}$ we get that, $\forall x_{1},x_{2}\in\mathbb{R}^{d}$
@@
from which follows
$\int\limits_{\mathbb{R}^{d}}g(\mathcal{L}(\pi,x))\mathrm{d}(t_{n\#}\pi-\pi)(x)-\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,x))\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z),t_{n}(x)-x\right\rangle\mathrm{d}\pi(x)$ (33)
$\geq-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}.$
Combining (30)–(33) we obtain
@@
$\;\phantom{\geq}-\int\limits_{\mathbb{R}^{d}}\int\limits_{\mathbb{R}^{d}}\left\langle g^{\prime}(\mathcal{L}(\pi,z))\xi_{2}(z,x),t_{n}(x)-x\right\rangle\mathrm{d}\pi(z)\mathrm{d}\pi(x)$
$\;\phantom{\geq}-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}.$